4 ms·
Because you're not guaranteed to roll a 6, even if you roll 6 times: 1/6 + 1/6 + 1/6 + 1/6 + 1/6 + 1/6 = 6/6 = 100% chance? Nope. :)
by 0x0 9y ago
Because you're not guaranteed to roll a 6, even if you roll 6 times: 1/6 + 1/6 + 1/6 + 1/6 + 1/6 + 1/6 = 6/6 = 100% chance? Nope. :)